5. As a US exporter, we have the option of hedging with forwards, options, or not hedging. The size of the transaction is 3.04 million Australian $. The current Exchange Rate is .7400-05 AUDUSD, and the 3 month forward Rate is .7500-05 AUDUSD. Australian Dollar options with an exercise price of .75 sells for a price of $.05 for calls and $.04 for puts. There are 100,000 Australian dollars in 1 option contract.

In 3 months when the transaction occurs, the final exchange rate is .700-02 AUDUSD, and option prices are .00 for calls and .10 for puts. (10)

a) How much would you receive/pay if you had hedged with forwards?

b) How much would you receive/pay if you had hedged with options?

c) How much would you receive/pay if you had not hedged?

d) Rank them
Answer #1

Let's start with the provided information

Given:

Current Exchange Rate: 0.74 Australian Dollar for every 5 US$.
[This means 1 USD= 6.76 AUD (5/0.74)]

3-month forward rate: 0.75 Australian Dollar for every 5 US$.
[This means 1 USD = 6.67 AUD (5/0.75), current view is bullish on Australian dollar, post 3 months]


Lot size for Options contract: 100000 Australian USD

However,
after 3 months the exchange rate is 0.70 Australian Dollar for every 2 US$.
[This means 1 USD= 2.86 AUD (2/0.70)]


This means the Australian currency is strengthening after 3 months. It means you can buy more US$ for less Australian Dollars. Although, exporters get paid in Home currency (here it is Australian Dollars). Strengthening of home currency is not favorable for exporters as now they get less Australian Dollars for every $ worth of goods sold. That is, today you are getting 6.76 AUD for every $, and post 3 months you will only get 2.86 AUD. So, loss of 3.9 AUD for every $ worth of goods post 3 months.

We can hedge a contract using Futures and options.

Futures is a forward contract that is either sold at discounted (less), premium (more) or at par (same) value as the spot (Current) rate. You can either short(Sell) or long (buy) the forward contract.

Options comprise of Calls (right to buy contract) and puts (right to sell contract) on the same spot price at the time of expiry. Calls below and equal to spot price are called "In the Money" (ITM) and puts above spot price are called Out of Money (OTM). As you are profited when you buy for less and sell for more, so you profit when the future price is more than spot price, if you are buyer. You buy call when you are bullish (anticipation that price will rise for undertaking).

Puts above and equal to spot price are called "In the Money" (ITM) and puts below spot price are called Out of Money (OTM). As you are profited when you sell for more today and buyback for less, so you profit when the future price is less than spot price if you are a seller. You buy puts when you are bearish (anticipation that price will fall for undertaking).

However, the future is uncertain and there is no certain way of knowing wheather price will rise or fall for the undertaking (here Australian Dollars)

a) How much would you receive/pay if you had hedged with forwards?

In Forwards contract, you can either long (buy) or short (sell) the 3-month forward contract, basis your anticipation of post 3-month price of Australian Dollars.

If you bearish, feel AUD price will fall (currency strengthening in this case) you can short (sell) the contract.
Sell 3-month forward rate: 0.75 Australian Dollar for every 5 US$.
[This means 1 USD = 6.67 AUD (5/0.75)]

Price you Pay here= Sell at 6.67,

post 3 months: the exchange rate is 0.70 Australian Dollar for every 2 US$.
[This means 1 USD= 2.86 AUD (2/0.70)]

Price you receive= Buyback at 2.86

Profit= 3.81 AUD

If you bullish, feel AUD price will rise (currency weakening in this case) you can long (buy) the contract.

Buy 3-month forward rate: 0.75 Australian Dollar for every 5 US$.
[This means 1 USD = 6.67 AUD (5/0.75)]

Price you Pay here= Buy at 6.67,

post 3 months: the exchange rate is 0.70 Australian Dollar for every 2 US$.
[This means 1 USD= 2.86 AUD (2/0.70)]

Price you receive= Sell at 2.86

Loss = 3.81 AUD

b) How much would you receive/pay if you had hedged with options?

OTM (out of money) Calls/Puts expire worthless at the time of expiry due to loss of time value of money.
Given:

In Options you have the option of buying/selling call or puts alone or use it in combination as a strategy.

If you are strongly bullish, you can buy Call options.
Amount you Pay= 0.05 *100000= 5000 AUD
Amount you receive= 0.00*100000= 0 AUD
Loss= 5000 AUD

If you are strongly bearish, you can buy Put options.
Amount you Pay= 0.04 *100000= 4000 AUD
Amount you receive= 0.10*100000= 10000 AUD
Profit= 6000 AUD

Or you can have a neutral strategy (neither bullish nor bearish)
You can buy a Straddle, i.e buy call and put of the same expiry and of same spot rate.
In this case,
Amount you Pay= (0.05 *100000) + (0.04 *100000)= 9000 AUD
Amount you receive= (0.00*100000) + (0.10*100000)= 10000 AUD
Profit= 6000 AUD

c) How much would you receive/pay if you had not hedged?

Not hedging is a high risk stategy and hedging protects you from sharp fluctuations in the currency.

If you dint hedge,
Current Exchange Rate: 0.74 Australian Dollar for every 5 US$.
[This means 1 USD= 6.76 AUD (5/0.74)]

Price you Pay here= Buy at 6.76,

post 3 months: the exchange rate is 0.70 Australian Dollar for every 2 US$.
[This means 1 USD= 2.86 AUD (2/0.70)]

Price you receive= Sell at 2.86

Loss = 3.9 AUD

d) Rank them

Ideally, a hedged contract is always preferred over the unhedged contract as hedging protects you from sharp fluctuations in price of undertaking.

Options give you the opportunity to have a neutral strategy, which is the case most of the time, as future price is always uncertain.

second best strategy is future contract, You can hedge yourself for future changes in price if you have a bullish or bearish view on the price of undertaking due to some information available to you.

Third and last option is unhedged strategy as it exposes you to unlimited loss/profit. In this case: Loss of 3.9 AUD